### v1

# Enhancing Retrosynthesis with Conformer: A Template-Free Method

Abstract
Retrosynthesis plays a crucial role in the fields of organic synthesis and drug development, where the goal is to identify suitable reactants that can yield a target product molecule. Although existing methods have achieved notable success, they typically overlook the 3D conformational details and internal spatial organization of molecules. This oversight makes it challenging to predict reactants that conform to genuine chemical principles, particularly when dealing with complex molecular structures, such as polycyclic and heteroaromatic compounds. In response to this challenge, we introduce a novel transformer-based, template-free approach that incorporates 3D conformer data and spatial information. Our approach includes an Atom-align Fusion module that integrates 3D positional data at the input stage, ensuring correct alignment between atom tokens and their respective 3D coordinates. Additionally, we propose a Distance-weighted Attention mechanism that refines the self-attention process, constricting the model’s focus to relevant atom pairs in 3D space. Extensive experiments on the USPTO-50K dataset demonstrate that our model outperforms previous template-free methods, setting a new benchmark for the field. A case study further highlights our method’s ability to predict reasonable and accurate reactants.
1 Introduction
Retrosynthesis, initially formulated by Corey (1991), stands as a fundamental problem within organic synthesis, aiming to discover and predict reactants given product molecules. This task presents a large chemical search space and myriad molecular combinations, posing significant challenges. Traditionally, researchers relied heavily on extensive knowledge of organic chemistry and reaction mechanisms, often resulting in inefficiency and heavy dependency on expertise. In recent years, with the significant impact of data-driven deep learning methods across various fields, computer-aided Retrosynthesis has gained increasing attention, especially the template-free methods.
Existing methods are categorized into three types based on reliance on template prior: template-based, semi-template, and template-free. Template-based methods Chen and Jung (2021); Coley et al. (2017); Dai et al. (2019) view Retrosynthesis as a template retrieval task and subsequently employ the retrieved templates (specific chemical changes) to transform the input product molecule into reactants using cheminformatics tools like RDKit Landrum and others (2013). Semi-template methods Shi et al. (2020); Yan et al. (2020) follow two steps: i) Reaction Center Identification: identifying the reaction centers of the product using product-reactant aligned atom mapping and then using RDKit to convert the product into synthons; ii) Synthons Completion: building another model to complete synthons into reactants. Template-free methods regard molecules as sequences, like the Simplified Molecular Input Line Entry System (SMILES) Weininger (1988), and frame Retrosynthesis prediction as machine translation. So that it can implicitly learn reaction rules and easily scale to larger data sets without relying on external template databases or molecular editing with RDKit.
Problem. In previous studies, the typical representation of molecules involves 1D sequences (i.e., SMILES) and 2D graphs. They overlooked the perception and comprehension of 3D conformer information, which is crucial for understanding molecular stereochemistry, reaction centers, chirality, and other properties Meng et al. (2023); Dang et al. (2018). Especially in molecules with intricate 3D structures (e.g., polychiral, heteroaromatic), the absence of 3D conformer information can lead to synthesis predictions that defy established chemical rules.
In this work, we focus on template-free Transformer for Retrosynthesis prediction. Incorporating 3D conformer information into this framework presents two challenges: i) integrating 3D features while maintaining alignment between atom tokens and corresponding 3D representations; ii) controlling receptive field of model based on spatial structure. To tackle these two challenges, we propose our method, a novel end-to-end Retrosynthesis Transformer that encodes both 1D SMILES sequences and 3D conformer information. We propose two modules, Atom-align Fusion and Distance-weighted Attention, to overcome the aforementioned challenges respectively. The model can jointly encode the 1D sequential and 3D positional information of molecules and redistribute attention weights based on 3D distances between atoms. Moreover, data augmentation Tetko et al. (2020); Zhong et al. (2022), like shuffling the atom order or aligning root atom between product and reactants, are employed to enhance the overall performance. Finally, we introduce SMILES alignment Deshpande and Narasimhan (2020); Wan et al. (2022) by incorporating attention guidance loss during training. Our model introduces 3D conformer information, allowing it to generate reasonable synthesis predictions even for molecules with intricate 3D structures, and it does not rely on additional molecular editing. We conducted experiments using a widely recognized dataset USPTO-50k Schneider et al. (2016). Results demonstrate that our model can enhance performance in terms of Top-k accuracy in both known and unknown reaction classes. Our proposed method achieves a new state-of-the-art for template-free methods, and it is also comparable to template-based and semi-template-based methods. Our contributions are summarized as follows:
  • To the best of our knowledge, the proposed model our method is the first to integrate 3D conformer information into template-free method and train the model in an end-to-end manner.
  • We propose i) Atom-align Fusion which integrates 1D and 3D representations while preserving their alignment. ii) Distance-weighted Attention which directs self-attention toward a more reasonable direction solely based on molecular 3D distance matrix.
  • Experiments on the USPTO-50k dataset indicate that proposed model our method achieves state-of-the-art Retrosynthesis performance in template-free methods with a Top-1 accuracy of 55.5% and 67.2% without and with reaction class, respectively.
2 Related Work
Template-based methods utilize reaction templates from pre-constructed database to capture the rules of chemical changes and match input products with the templates. RetroSim Coley et al. (2017) is a similarity-based approach using molecular fingerprint similarity to rank candidate templates. GLN Dai et al. (2019) employs the Conditional Graph Logic Network to learn when rules from reaction templates should be applied, implicitly considering chemical feasibility and strategy. Inspired by the chemical intuition that molecular changes primarily occur locally during reactions, LocalRetro Chen and Jung (2021) is locally encoded and refined to accommodate non-local reactions. Although these methods perform the best among the three categories, their reliance on pre-defined templates limits their coverage and scalability, rendering them insufficient for real-world scenarios.
Semi-template methods do not directly use templates but instead combine chemical rules with generative models. Existing methods typically follow a two-stage procedure: i) reaction center identification; ii) synthons completion. Reaction center identification predicts the reaction center of the product and breaks the corresponding bond into incomplete molecules, called synthons. Synthons completion uses generative models to complete synthons into reactants. RetroXpert Yan et al. (2020) builds an Edge-enhanced Graph Attention Network (EGAT) that takes a molecule graph as input and predicts bond disconnection probabilities to obtain synthons. It then employs a seq2seq model to generate reactants in SMILES format. In contrast to the SMILES format, G2Gs Shi et al. (2020) completes synthons into reactants in graph format through node selection and edge labeling. GraphRetro Somnath et al. (2020) expands synthons into complete reactants by attaching relevant leaving groups. RetroPrime Wang et al. (2021) introduces a two-stage procedure by adding additional labels to SMILES. In contrast, MEGAN Nijssen and Kok (2004) and Graph2Edits Zhong et al. (2023) reframe the generative process as a series of graph edits executed by RDKit.
[…]
Figure 1: Architecture overview of our method based on Transformer. a) SMILES and molecular conformer are processed through Word Embedding and ComENet to yield SMILES sequence embedding and 3D position embedding. Following the Atom-align integration, fusion embedding is then input into the Encoder. b) In Encoder, 3D Distance weight is obtained for redistribution of self-attention by Gaussian Basis Kernel function, while also refined layer-by-layer. c) In Decoder, SMILES Alignment is employed as guidance to align the cross-attention between the product and reactants.
3 Preliminary
3.1 Template-free Retrosynthesis
Molecules can be represented as sequences of tokens in SMILES format. Following the SMILES tokenization of Schwaller et al. (2019) that separates each atom (i.e., O, C, N, Cl) and non-atom tokens (i.e., -, =, #), we let $S=[s_{1},s_{2},\dots s_{n}]$ denote SMILES sequence with $n$ number of tokens. In template-free Retrosynthesis, we obtain the product SMILES $S_{P}$ and the reactants SMILES $S_{R}$, and attempt to find a method $f_{\theta}$ that can map $S_{P}$ to $S_{R}$ with auto-regressive way: $S_{R}=f_{\theta}(s_{i+1}|s_{0},\dots,s_{i},S_{P})$ in $i$ step.
3.2 Molecular Conformer and ComENet
Molecule is a dynamic structure since atoms are in continual motion in 3D space. The local minima on the potential energy surface are called conformers. Generally, molecular conformer can be represented as $G_{\text{3D}}=(A,C)$, where $A=\{a_{i}\}_{i=1,\dots,n}$ is atoms list, $C=\{(x_{i},y_{i},z_{i})\}_{i=1,\dots,n}$ is a 3D coordinate list. To fulfill the global completeness of 3D conformer information of a molecule, ComENet Wang et al. (2022) encodes rotation angles and spherical coordinates within the 1-hop neighborhood of atoms into the message passing scheme Balcilar et al. (2021).
4 Method
In this section, we introduce our method, a novel transformer-based model that is capable of learning richer molecular representation through SMILES sequence and spatial conformer. An overview of the proposed method is shown in Figure 1.
Figure 1a illustrates Atom-align Fusion that incorporates both 1D SMILES tokens embedding and 3D position embedding to get fusion embedding as inputs for encoder-decoder model. By adopting this approach, the model is capable of integrating 3D positional information while preserving alignment (Section 4.1). Figure 1b illustrates Distance-weighted Attention, which directs self-attention to prioritize chemical relevant atoms within spatial structure (Section 4.2). Figure 1c illustrates SMILES alignment strategy for further bridging the mapping relationship between product SMILES and reactant SMILES in decoder cross-attention (Section 4.3). The overall training and inference process is conducted in an end-to-end manner, and it is a template-free method that does not rely on RDKit for molecular editing.
4.1 Atom-align Fusion
The most direct method of feature integration is concatenation, but this approach disrupts the alignment between 1D representations of atom tokens and their corresponding 3D representations. The challenge lies in how to integrate conformer information into the molecular representation while preserving alignment between them.
To address this issue, we introduced the Atom-align Fusion module. The purpose is to obtain fusion embedding $F_{\text{3D}}\in\mathbb{R}^{M\times D}$ by associating SMILES token embedding $T\in\mathbb{R}^{M\times D}$ with 3D position embedding $P_{\text{3D}}\in\mathbb{R}^{N\times D}$, where $D$ is model dimension, $M$ is number of SMILES tokens, $N$ is number of atoms.
Firstly, to obtain the complete 3D molecular representation, we extracted 3D position embedding $P_{\text{3D}}$ from the molecular conformer using ComENet Wang et al. (2022). ComENet proposes a novel message passing scheme Balcilar et al. (2021) that operates within a 1-hop neighborhood with radial distance $d$, polar angle $\theta$, azimuthal angle $\phi$ and rotation angle $\tau$, and then builds message passing scheme as
$\mathbf{v}_{i}^{t+1}=g\left(\mathbf{v}_{i}^{t},\sum_{j\in\mathcal{N}_{i}}f\left(\mathbf{v}_{j},d_{ij},\theta_{ij},\phi_{ij},\tau_{ij}\right)\right)$ (1)
where $\mathbf{v}_{i}^{t}\rightarrow\mathbf{v}_{i}^{t+1}$ means update of atom (node) representation, $g$ and $f$ implemented by neural networks or mathematical operations, $d_{ij}$, $\theta_{ij}$ and $\phi_{ij}$ specify the 1-hop local neighborhood, $\tau_{ij}$ determines the orientation of the local neighborhood. We omit the final aggregation layer of ComENet to obtain the 3D representation corresponding to each atom $P_{\text{3D}}=\{\mathbf{v}_{i}\}_{i=1,2,\dots,N}$ rather than the entire molecular graph.
Subsequently, we input $P_{\text{3D}}$ along with the SMILES tokens embedding $T$ into the encoder. To facilitate a more comprehensive integration of SMILES and conformer, we aligned tokens embedding $T$ and $P_{\text{3D}}$ with atom index in the same vector space. We initially enlarge $P_{\text{3D}}$ from $N$ (number of atoms) length to $M$ (number of tokens) length by padding 0 to the indexes of non-atom tokens. In model input stage, we blend the $P_{\text{3D}}$ and $T$ to get the fusion embedding $F_{\text{3D}}$, allowing effective integration of 3D positional and sequential information at each atom position while preserving the alignment between them:
$F_{\text{3D}}=\lambda_{1}\times\operatorname{pad}(P_{\text{3D}})+\lambda_{2}\times T$ (2)
where $\lambda_{1}$ and $\lambda_{2}$ are trainable parameters that can adaptively adjust weights of 3D position embedding $P_{\text{3D}}$ and SMILES tokens embedding $T$ during training. Besides, we still use Sinusoidal Position Embedding Vaswani et al. (2017) for tagging position of tokens.
4.2 Distance-weighted Attention
According to Li et al. (2017); Wan et al. (2022), multi-head attention calculation in vanilla transformer is insufficient and lacks chemical knowledge. Simultaneously, 3D distances between atoms within a molecule reflect the spatial relation between any pair of atoms in physical chemistry, such as bond energy, stability, electron affinity and interatomic force. Incorporating 3D distances into the multi-head attention calculation process enables the model to reschedule attention accordingly, which not only alleviates the redundancy in multi-head attention but also enables distance-based attention redistribution during the attention calculation.
Similar to the method of partitioning attention heads in Wan et al. (2022), we divide the multi-head attention into normal attention heads and spatial attention heads. Normal attention heads follow the traditional transformer manner to learn sequence-wide information, whereas spatial attention heads employ 3D distances as an external attention weight to enable attention redistributed based on 3D distances. The whole module can be formulated as two steps: 1) Construct 3D distance weight, 2) Attention Redistribute & Weight Refine.
Construct 3D Distance Weight
First, we construct a distance matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ from molecular conformer of molecule as
$\mathbf{D}=\{\mathbf{d}_{i,j}\}_{i,j=1,2,\dots,N}\quad\mathbf{d}_{ij}=\|\mathbf{c}_{i}-\mathbf{c}_{j}\|_{2}$ (3)
where $N$ is number of atoms, $i,j$ are index of atoms pair, $\mathbf{c}_{i}$ is coordiante of $i$-th atom. To explore high-dimensional space, we employ Gaussian Basis Function to transform the distance $\mathbf{d}_{ij}$ between each atom pair ($i$, $j$) into a higher dimension as
$\psi_{(i,j)}^{k}=-\frac{1}{\sqrt{2\pi}\left|\sigma^{k}\right|}\exp\left(-\frac{1}{2}\left(\frac{\gamma_{(i,j)}\mathbf{d}_{ij}+\beta_{(i,j)}-\mu^{k}}{\left|\sigma^{k}\right|}\right)^{2}\right)$
where k={1, …, K}, K is the number of Gaussian Kernel kernels. $\mu^{k}$ and $\sigma^{k}$ represent kernel center and scaling factor of $k$-th Gaussian Basis kernel. $\gamma_{(i,j)}$ and $\beta_{(i,j)}$ are learnable scalars related to bond type (i.e., SINGLE, DOUBLE) of pair ($i$, $j$). To obtain the 3D Distance weight $\Phi_{(i,j)}$, we use two layers of Non-Linear Layer with the activation function Gaussian Error Linear Units ($\operatorname{GELU}$) Hendrycks and Gimpel (2016) to encode $\psi_{(i,j)}$ as
$\Phi_{(i,j)}=\mathbf{W_{2}}(\operatorname{GELU}(\mathbf{W_{1}}(\psi_{(i,j)})))$ (5)
where $\psi_{(i,j)}=\operatorname{stack}([\psi_{(i,j)}^{1};\dots;\psi_{(i,j)}^{k}])$, $\mathbf{W_{1}},\mathbf{W_{2}}\in\mathbb{R}^{K\times K}$ are learnable parameters and 3D Distance weight $\Phi_{(i,j)}\in\mathbb{R}^{N\times N\times K}$.
Attention Redistribute & Weight Refine
Second, we leverage the previously obtained 3D Distance weight $\psi_{(i,j)}$ to guide the model in redistributing attention by the internal spatial structure, instead of directly calculating attention scores between each token like vanilla Transformer. We divide the attention heads into two halves: normal attention head and spatial attention head. The normal attention head still captures the context of the SMILES sequence like original attention computation. The spatial attention head, on the other hand, focuses more on the molecular internal 3D structure, limiting the receptive field of each atom and considering the 3D distance with other atoms. For the $i$-th token of the $l$-th encoder layer, the roll-out form of spatial attention head is formulated as:
$x_{i,\text{spatial}}^{l+1}$ $=\sum_{j\in N(i)}\operatorname{softmax}\left(\Phi_{(i,j)}^{l}\odot\frac{q_{i}k_{j}^{T}}{\sqrt{d}}\right)v_{j}$ (6)
${\left[q_{i},k_{j},v_{j}\right]}$ $=\left[h_{i}^{l}\mathbf{W^{Q}},h_{j}^{l}\mathbf{W^{K}},h_{j}^{l}\mathbf{W^{V}}\right]$
where $\Phi_{(i,j)}^{l}$ is almost the same as 3D Distance weight $\Phi_{(i,j)}$ but padding 0 in non-atom token for element-wise multiplication $\odot$, $\mathbf{W^{Q}}$, $\mathbf{W^{K}}$ and $\mathbf{W^{V}}$ are projection matrices for query $q$, key $k$, value $v$ in self-attention. Computed representations of normal attention head and spatial attention head are concatenated along the hidden dimension and then processed through a linear layer, forming the updated context for the next layer $h^{l+1}$ as:
$h^{l+1}=\mathbf{W_{Linear}}\left(\left[x^{l+1}_{\text{normal}};x^{l+1}_{\text{spatial}}\right]\right)$ (7)
In addition, we perform Weight Refine to update our 3D Distance weight layer by layer. The Weight Refine module is a fully connected (FC) layer that takes concatenation of updated contexts of atoms pair ($i$, $j$) to refine the 3D attention weight among receptive field:
$\Phi_{(i,j)}^{l+1}=\Phi_{(i,j)}^{l}+\operatorname{FC}\left(\left[h_{i}^{l+1};h_{j}^{l+1}\right]\right)$ (8)
4.3 Training Strategy
SMILES Alignment
During the occurrence of chemical reactions, the majority of atoms remain unchanged. We can utilize the atom-map to annotate the correspondence between atoms of product and reactants in reaction, which remains unchanged during the reaction. Furthermore, the correspondence of atoms can be readily transferred to the token correspondence in the SMILES sequence (see in Figure 1c) and construct SMILES Alignment Map (SAM) as:
[…]
Inspired by Deshpande and Narasimhan (2020), we similarly employed the guided attention approach by introducing guidance loss $\mathcal{L}_{SA}$ to encourage alignment between the cross attention scores in the final layer of the decoder and SMILES Alignment Map, strengthening the attention weights between corresponding tokens.
Data Augmentation
When molecules are converted into SMILES, different choices of root atoms can result in different SMILES representations, and this non-uniqueness may impede training of seq2seq models. Apart from unifying SMILES into a canonical form, another approach is to transform and enlarge the original SMILES sets through data augmentation. Drawing from some effective data augmentation tricks in prior research Dai et al. (2019); Tetko et al. (2020); Zhong et al. (2022), we explore two types of data augmentation approaches: Randomize SMILES and Root-align SMILES.
  • Randomize SMILES: By shuffling the order of atoms in a molecule or altering the priority of side chains, different SMILES representations can be generated. There can be up to n! (where n is the number of atoms in the molecule) possible SMILES to represent a molecule.
  • Root-align SMILES (R- SMILES): First, a root atom is randomly selected from product, and then product R-SMILES is generated based on this root atom. Next, the same root atom in reactants is identified based on atom mapping or substructure matching between product and reactants and generate reactants R-SMILES. R-SMILES effectively reduces the editing distance between product and reactants, ensuring a high sequence similarity and simplifying the prediction problem.
    R-product: Brc1ccc(C=O)cn1
    R-reactants: Brc1ccc(Br)cn1.C(N(C)C)=O
In contrast to other methods that directly expand dataset, we employ an on-the-fly approach to progressively augment data during training iterations. The specific procedure involves transforming the input canonical SMILES into Randomize SMILES (or Root-align SMILES) with a 50% probability in each training iteration. As iterations increase, the model gradually gains a broader perspective on permuted SMILES. This dynamic augmentation approach allows the model to focus more on the standard form (canonical SMILES) while regularizing with the permuted form (permuted SMILES), thereby improving the accuracy, stability, and generalization of model.
     Model Top-k accuracy (%)
Reaction class known Reaction class unknown
1 3 5 10 1 3 5 10
Template-Based
RetroSim 52.9 73.8 81.2 88.1 37.3 54.7 63.3 74.1
GLN 64.2 79.1 85.2 90.0 52.5 69.0 75.6 83.7
LocalRetro 63.9 86.8 92.4 96.3 53.4 77.5 85.9 92.4
Semi-Template-Based
RetroXpert 62.1 75.8 78.5 80.9 50.4 61.1 62.3 63.4
G2G 61.0 81.3 86.0 88.7 48.9 67.6 72.5 75.5
GraphRetro 63.9 81.5 85.2 88.1 53.7 68.3 72.2 75.5
RetroPrime 64.8 81.6 85.0 86.9 51.4 70.8 74.0 76.1
MEGAN 60.7 82.0 87.5 91.6 48.1 70.7 78.4 86.1
Graph2Edits 67.1 87.5 91.5 93.6 55.1 77.3 83.4 89.4
Template-Free
SCROP 59.0 74.8 78.1 81.1 43.7 60.0 65.2 68.7
GET 57.4 71.3 74.8 77.4 44.9 58.8 62.4 65.9
Aug. Transformer - - - - 48.3 - 73.4 77.4
GTA - - - - 51.1 67.6 74.8 81.6
Graph2SMILES - - - - 52.9 66.5 70.0 72.9
Retroformer 64.0 82.5 86.7 90.2 53.2 71.1 76.6 82.1
our method (Ours) 65.3 84.0 88.3 91.6 53.6 73.7 80.2 86.1
our method (Ours, +R) 67.2 87.0 91.4 94.4 55.5 76.3 82.8 88.8
Table 1: Top-k accuracy for Retrosynthesis prediction on USPTO-50K. +R indicates R-SMILES augmentation. The metric values of other models in the table are taken from their original papers. Best performance is in bold.
     Model Top-k validity (%)
1 3 5 10
Graph2SMILES 99.4 90.9 84.9 74.9
RetroPrime 98.9 98.2 97.1 92.5
Retroformer 99.2 98.5 97.4 96.7
our method (Ours) 99.8 99.6 99.1 97.0
our method (Ours, +R) 99.8 99.4 98.6 95.4
Table 2: Top-k SMILES validity for Retrosynthesis prediction on USPTO-50K with reaction class unknown.
4.4 Overall Loss
The overall loss is composed of the language modeling R-Drop lossWu et al. (2021) $\mathcal{L}_{LM}$ for Retrosynthesis, along with the SMILES Alignment guidance loss $\mathcal{L}_{SA}$. R-drop Loss builds upon cross-entropy loss enhancing prediction consistency and model robustness by minimizing bidirectional KL-divergence of output distributions of any pair of sub-models sampled from dropout in model training.
The objectives of overall loss: i) approximate target in retrosynthesis, ii) ensure the robustness under different dropout, iii) strengthen the alignment relationship in cross-attention. The formula is as follows:
$\mathcal{L}=\mathcal{L}_{LM}^{(CE)}+\alpha\mathcal{L}_{LM}^{(KL)}+\mathcal{L}_{SA}$ (10)
where $\mathcal{L}_{LM}^{(CE)}$ is Cross-Entropy loss to minimize prediction error against actual outputs, $\mathcal{L}_{LM}^{(KL)}$ is KL-divergence in R-Drop loss to strive for outputs of models with varying Dropout to be as similar as possible, and we set $\alpha$ to 0.5. $\mathcal{L}_{SA}$ is Cross-Entropy loss between cross-attention score of last decoder layer and SMILES Alignment Map that we mention in Section 4.3.
5 Experiments
Datasets
We use the popular Retrosynthesis benchmark dataset USPTO-50K Schneider et al. (2016), which contains 50,016 atom-mapped reactions, to evaluate our model. The same data split as Coley et al. (2017) is applied to our experiments, resulting in 40,008, 5,001, and 5,007 reactions for the training, validation, and test sets, respectively. Conformers are generated using RDKit with the stochastic optimization algorithm Merck Molecular Force Field (MMFF). We then apply the same algorithm in Wan et al. (2022) to extract the ground truth SMILES Alignment Map.
Setting
Our model is built on vanilla Transformer, consisting of 8 encoder layers and 8 decoder layers with 8 attention heads. The model dimension is set to 512. We employ the Adam optimizer Kingma and Ba (2014) with $(\beta_{1},\beta_{2})=(0.9,0.98)$ and train the model with batch size of 16. During training, we set an early-stop of 7 epochs under a maximum of 1000 epochs to obtain the 7 best models and average their parameters. Finally, the whole training process took approximately 30 hours on a single NVIDIA GeForce RTX 4090 GPU.
Evaluation
During inference, we use beam search Tillmann and Ney (2003) decode strategy with a beam size is 10. After that, we evaluate our model with Top-k accuracy, which considers it correct only if Top-k predicted reactants of the beam search results are the same as original test set. For molecule validity, we verify Top-k validity based on whether RDKit can recognize the Top-k predicted reactants.
Baseline
We compare our proposed method, our method, with several strong baseline models from three types of Retrosynthesis method. We take RetroSim Coley et al. (2017), GLN Dai et al. (2019) and LocalRetro Chen and Jung (2021) to represent template-based methods. We take RetroXpert Yan et al. (2020), G2Gs Shi et al. (2020), GraphRetro Somnath et al. (2020), RetroPrime Wang et al. (2021), MEGAN Nijssen and Kok (2004) and Graph2Edits Zhong et al. (2023) to represent semi-template method. We take SCROP Zheng et al. (2019), Aug.Transformer Tetko et al. (2020), Graph2SMILES Zhong et al. (2023), GET Mao et al. (2021), GTA Seo et al. (2021) and Retroformer Wan et al. (2022) to represent template-free method. We did not choose pre-training methods for comparison as they are trained on larger datasets. We experimented with two training strategies for our proposed model: our method represents the model with Randomize SMILES data augmentation; our method+R represents the model with Root-align SMILES data augmentation.
(a) Polychiral
(b) Heteroaromatic
(c) Fused / Bridged Rings
(d) Complex
Figure 2: Case study for different structures. Each subfigure shows predictions of different methods: our method (Top), Transformer (Bottom).
5.1 Performance
Top-k Accuracy
As shown in Table 1, for the model with random augmentation our method, when reaction class is known, Top-1/Top-10 accuracy is 65.3%/91.6%, exceeding previous state-of-the-art template-free methods. When reaction class is unknown, our model achieves Top-1/Top-10 accuracy at 53.6%/86.1% and also outperforms existing template-free methods. It’s worth noting that model demonstrates competitive performance in both template-based and semi-template methods, even though they rely on predefined templates or RDKit editing. For the model with R-SMILES augmentation our method+R, it is astonishingly found that by further aligning the root of product and reactants, our model not only greatly surpasses state-of-art template-free methods, but also essentially exceeds other template-based methods and semi-template methods on several metrics.
Top-k Validity
We take three typical template-free methods Graph2SMILES, RetroPrime and Retroformer as baseline models to verify the predicted reactants SMILES validity of our model. We refrained from selecting template-based and semi-template methods, as SMILES constructed based on template information ensure validity. Conversely, template-free methods often struggle to generate valid SMILES without extra chemical guidance. As shown in Table 2, our models are more capable of generating valid SMILES than other baseline models. Especially Top5-10 reactants, which are far ahead of other models. This indicates that incorporating 3D conformer information can assist the model in grasping the structures involved in chemical reactions, thus preventing the generation of SMILES that contravene chemical principles.
  (a) (b) Top-k accuracy (%)
1 3 5 10
- - 50.9 71.2 78.1 83.7
$\checkmark$ - 51.4 70.5 76.5 81.9
- $\checkmark$ 52.8 72.1 78.2 84.0
$\checkmark$ $\checkmark$ 53.6 73.7 80.2 86.1
Table 3: Ablation study of our method with reaction class unknown and random data augmentation. Module (a) is Atom-align Fusion; Module (b) is Distance-weighted Attention;
5.2 Ablation Study
To explore the impact of two modules, we design an ablation study under the settings of unknown reaction class and random SMILES data augmentation. The results are shown in Table 3 in the form of top-k accuracy and demonstrate necessity of each module for our method. The improvement in top-1 accuracy from ② to ① suggests that the model is able to use 3D positional information to infer the correct results with higher confidence after introducing module (a) Atom-align Fusion. Notably, the declines in top 3-10 accuracy is speculated on the reliance upon fixed 3D position embedding, which constrains the search space of model. Comparing ③ and ①, it can be seen that employing (b) Distance-weighted Attention improves performance in all top-k accuracy, even for top 3-10. We conjecture that redistributing attention enables the model to focus on internal structure of molecules, while Gaussian Kernel Function grants the model a larger search space. The results of ④ show that the combination of (a) and (b) has achieved the best performance of model, enabling it to focus on internal spatial structures while maintaining an extensive search space based on comprehensive 3D positional information.
5.3 Case Study
To verify that the proposed method our method can handle molecules with more intricate spatial structures, we chose 4 types of products as representatives:
  • Polychiral: a product with two or more chiral centers (’@/@@’ in SMILES), leading to complex stereochemistry and multiple isomers.
  • Heteroaromatic: a product containing at least one heteroatom as part of its aromatic ring structure.
  • Fused / Bridged Rings: product features multiple ring structures joined together, either sharing common atoms or connected by bridge atoms.
  • Complex: a product encompassing two or more features mentioned above.
Regarding the 4 types of products, we experiment on our method and a Transformer without 3D information and choose the Top1-3 predictions for analysis, as shown in Figure 2. The rows of each subfigure correspond to our method and Transformer, respectively. Ground Truth and invalid SMILES are both highlighted in bold red.
The results show that when dealing with molecules of intricate spatial structures, our method can infer reasonable reactants with high confidence (often at top1, top2), while Transformer without 3D information finds it challenging to ascertain accurate results. Particularly in the case of Complex products, three predictions of our method could deduce skeleton and achieve the ground truth at top-1, whereas Transformer could not get any valid SMILES at top1-3. This further confirms the importance of introducing 3D information when dealing with molecules of intricate spatial structures.
6 Conclusion
We propose our method, an end-to-end Transformer that integrates 3D conformer for Retrosynthesis. With the proposed Atom-align Fusion module, we can adaptively integrate token and 3D position information ensuring the alignment between them. Next, we propose Distance-weighted Attention mechanism to guide the redistribution of self-attention with spatial distances and refine it across multiple layers. Our model achieves a new state-of-the-art in template-free methods and is also highly competitive with template-based and semi-template methods. In future work, we will attempt to introduce our method into multi-step synthesis route planning.