### v2

# Retro3D: A 3D-aware Template-free Method for Enhancing Retrosynthesis via Molecular Conformer Information

Abstract
Retrosynthesis plays a crucial role in organic synthesis and drug discovery, focusing on identifying a set of reactants capable of synthesizing a target product molecule. Although the existing approaches have shown promising results, they often overlook 3D conformer information and molecular spatial structure, limiting their ability to generate reactants that comply with chemical rules, especially for complex molecules (e.g., polychiral, heteroaromatic). To tackle this problem, we propose Retro3D, a transformer-based template-free method that integrates molecular conformer information and spatial structure. We devise an Atom-align Fusion module to combine 3D positional information in model input stage, ensuring alignment between atom tokens and corresponding 3D representations. Furthermore, we design a Distance-weighted Attention mechanism to guide self-attention, constraining receptive field of model and emphasizing chemical relevant atom pairs in 3D space. Experiments conducted on the USPTO-50K and USPTO-FULL datasets demonstrate that Retro3D significantly outperforms existing template-free approaches, achieving a new state-of-the-art performance. Case studies further highlight its capability to predict accurate and chemically plausible reactants, even for target molecules with intricate structures.
1 Introduction
Retrosynthesis, initially formulated by Corey (1991), stands as a fundamental problem within organic synthesis, aiming to discover and predict reactants given product molecules. This task presents a large chemical search space and myriad molecular combinations, posing significant challenges. Traditionally, researchers relied heavily on extensive knowledge of organic chemistry and reaction mechanisms, often resulting in inefficiency and heavy dependency on expertise. In recent years, with the significant impact of data-driven deep learning methods across various fields, computer-aided Retrosynthesis has gained increasing attention, especially the template-free methods Zhong et al. (2024).
Existing methods are categorized into three types based on reliance on template prior: template-based, semi-template, and template-free. Template-based methods Chen and Jung (2021); Coley et al. (2017); Dai et al. (2019) view Retrosynthesis as a template retrieval task and subsequently employ the retrieved templates (specific chemical changes) to transform the input product molecule into reactants using cheminformatics tools like RDKit Landrum and others (2013). Semi-template methods Shi et al. (2020); Yan et al. (2020) follow two steps: i) Reaction Center Identification: identifying the reaction centers of the product using product-reactant aligned atom mapping and then using RDKit to convert the product into synthons; ii) Synthons Completion: building another model to complete synthons into reactants. Template-free methods Zheng et al. (2019); Tetko et al. (2020); Mao et al. (2021) regard molecules as sequences, like the Simplified Molecular Input Line Entry System (SMILES) Weininger (1988), and frame Retrosynthesis prediction as machine translation. So that it can implicitly learn reaction rules and easily scale to larger data sets without relying on external template databases or molecular editing with RDKit.
Problem. In previous studies, the typical representation of molecules involves 1D sequences (i.e., SMILES) and 2D graphs. They overlooked the perception and comprehension of 3D conformer information, which is crucial for understanding molecular stereochemistry, reaction centers, chirality, and other properties Meng et al. (2023); Dang et al. (2018). Especially in molecules with intricate 3D structures (e.g., polychiral, heteroaromatic), the absence of 3D conformer information can lead to synthesis predictions that defy established chemical rules Imrie et al. (2020).
In this work, we focus on template-free Transformer for Retrosynthesis prediction. Incorporating 3D conformer information into this framework presents two challenges: i) integrating 3D features while maintaining alignment between atom tokens and corresponding 3D representations; ii) constraining receptive field of model based on spatial structure. To tackle these two challenges, we propose Retro3D, a novel end-to-end Retrosynthesis Transformer that encodes both 1D SMILES sequences and 3D conformer information. We propose two modules, Atom-align Fusion and Distance-weighted Attention, to overcome the aforementioned challenges respectively. The model can jointly encode the 1D sequential and 3D positional information of molecules and redistribute attention weights based on 3D distances between atoms. Moreover, data augmentation Tetko et al. (2020); Zhong et al. (2022), like shuffling the atom order or aligning root atom between product and reactants, are employed to enhance the overall performance. Finally, we introduce SMILES alignment Deshpande and Narasimhan (2020); Wan et al. (2022) by incorporating attention guidance loss during training. Our model introduces 3D conformer information, allowing it to generate reasonable synthesis predictions even for molecules with intricate 3D structures, and it does not rely on additional molecular editing. We conducted experiments using two widely recognized datasets: USPTO-50k Schneider et al. (2016), USPTO-FULL. Results demonstrate that our model can enhance performance in terms of Top-k accuracy in both known and unknown reaction classes. Our proposed method achieves a new state-of-the-art for template-free methods, and it is also comparable to template-based and semi-template-based methods. Our contributions are summarized as follows:
  • We propose Retro3D, an end-to-end trained, template-free method that integrates 3D conformer information to generate synthesis routes more in line with real-world scenarios for intricate molecular structures.
  • We propose i) Atom-align Fusion which integrates 1D and 3D representations while preserving their alignment. ii) Distance-weighted Attention which directs self-attention toward a more reasonable direction solely based on molecular 3D distance matrix.
  • Experimental results on the USPTO-50k and USPTO-FULL datasets demonstrate that the proposed model, Retro3D, achieves state-of-the-art retrosynthesis performance among template-free methods, under both with and without reaction class conditions.
2 Related Work
Template-based methods utilize reaction templates from pre-constructed database to capture the rules of chemical changes and match input products with the templates. RetroSim Coley et al. (2017) is a similarity-based approach using molecular fingerprint similarity to rank candidate templates. GLN Dai et al. (2019) employs the Conditional Graph Logic Network to learn when rules from reaction templates should be applied, implicitly considering chemical feasibility and strategy. Inspired by the chemical intuition that molecular changes primarily occur locally during reactions, LocalRetro Chen and Jung (2021) is locally encoded and refined to accommodate non-local reactions. Although these methods have strong interpretability, their reliance on pre-defined templates limits their coverage and scalability, rendering them insufficient for real-world scenarios.
Semi-template methods do not directly use templates but instead combine chemical rules with generative models. Existing methods typically follow a two-stage procedure: i) reaction center identification; ii) synthons completion. Reaction center identification predicts the reaction center of the product and breaks the corresponding bond into incomplete molecules, called synthons. Synthons completion uses generative models to complete synthons into reactants. RetroXpert Yan et al. (2020) builds an Edge-enhanced Graph Attention Network (EGAT) that takes a molecule graph as input and predicts bond disconnection probabilities to obtain synthons. It then employs a seq2seq model to generate reactants SMILES. In contrast to SMILES format, G2Gs Shi et al. (2020) completes synthons into reactants in graph format through node selection and edge labeling. GraphRetro Somnath et al. (2020) expands synthons into complete reactants by attaching leaving groups. RetroPrime Wang et al. (2021) introduces a two-stage procedure by adding additional labels to SMILES. In contrast, MEGAN Sacha et al. (2021) and Graph2Edits Zhong et al. (2023) reframe the generative process as a series of graph edits executed by RDKit.
[…]
3 Preliminary
3.1 Template-free Retrosynthesis
Molecules can be represented as sequences of tokens in SMILES format. Following the SMILES tokenization of Schwaller et al. (2019) that separates each atom (i.e., O, C, N, Cl) and non-atom tokens (i.e., -, =, #), we let $S=[s_{1},s_{2},\dots s_{n}]$ denote SMILES sequence with $n$ tokens. In template-free Retrosynthesis, we obtain the product SMILES $S_{P}$ and the reactants SMILES $S_{R}$, and attempt to find a method $f_{\theta}$ that can map $S_{P}$ to $S_{R}$ in auto-regressive way: $S_{R}=f_{\theta}(s_{i+1}|s_{0},\dots,s_{i},S_{P})$ in $i$ step.
3.2 Molecular Conformer and ComENet
Molecule is a dynamic structure since atoms are in continual motion in 3D space. The local minima on the potential energy surface are called conformers. Generally, molecular conformer can be represented as $G_{\text{3D}}=(A,C)$, where $A=\{a_{i}\}_{i=1,\dots,n}$ is atoms list, $C=\{(x_{i},y_{i},z_{i})\}_{i=1,\dots,n}$ is 3D coordinates list. To fulfill the global completeness of 3D conformer information of a molecule, ComENet Wang et al. (2022) encodes rotation angles and spherical coordinates within the 1-hop neighborhood of atoms via the message passing scheme Balcilar et al. (2021).
Figure 1: Architecture overview of Retro3D based on Transformer. a) SMILES and molecular conformer are processed through Word Embedding and ComENet to yield SMILES sequence embedding and 3D position embedding. Following the Atom-align integration, fusion embedding is then input into the Encoder. b) In Encoder, 3D Distance weight is obtained for redistribution of self-attention by Gaussian Basis Kernel function, while also refined layer-by-layer. c) In Decoder, SMILES Alignment is employed as guidance to align the cross-attention between the product and reactants.
4 Method
In this section, we introduce Retro3D, a novel transformer-based model that is capable of learning richer molecular representation through SMILES sequence and spatial conformer. An overview of the proposed method is shown in Figure 1.
Figure 1a illustrates Atom-align Fusion that incorporates both 1D SMILES tokens embedding and 3D position embedding to get fusion embedding as inputs for encoder-decoder model. By adopting this approach, the model is capable of integrating 3D positional information while preserving alignment (Section 4.1). Figure 1b illustrates Distance-weighted Attention, which directs self-attention to prioritize chemical relevant atoms within spatial structure (Section 4.2). Figure 1c illustrates SMILES alignment strategy for further bridging the mapping relationship between product SMILES and reactant SMILES in cross-attention module of decoder (Section 4.3). The overall training and inference process is conducted in an end-to-end manner, and it is a template-free method that does not rely on RDKit for molecular editing.
4.1 Atom-align Fusion
The most direct method of feature integration is concatenation Mao et al. (2021), but this approach disrupts the alignment between 1D representations of atom tokens and their corresponding 3D representations. The challenge lies in how to integrate conformer information into the molecular representation while preserving alignment between them.
To address this issue, we introduced the Atom-align Fusion module. The purpose is to obtain fusion embedding $F_{\text{3D}}\in\mathbb{R}^{M\times D}$ by associating SMILES token embedding $T\in\mathbb{R}^{M\times D}$ with 3D position embedding $P_{\text{3D}}\in\mathbb{R}^{N\times D}$, where $D$ is model dimension, $M$ is number of SMILES tokens, $N$ is number of atoms.
Firstly, to obtain the complete 3D molecular representation, we extracted 3D position embedding $P_{\text{3D}}$ from the molecular conformer using ComENet Wang et al. (2022). ComENet proposes a novel message passing scheme Balcilar et al. (2021) that operates within a 1-hop neighborhood with radial distance $d$, polar angle $\theta$, azimuthal angle $\phi$ and rotation angle $\tau$, and then builds message passing scheme as:
$\mathbf{v}_{i}^{t+1}=g\left(\mathbf{v}_{i}^{t},\sum_{j\in\mathcal{N}_{i}}f\left(\mathbf{v}_{j},d_{ij},\theta_{ij},\phi_{ij},\tau_{ij}\right)\right),$ (1)
where $\mathbf{v}_{i}^{t}\rightarrow\mathbf{v}_{i}^{t+1}$ means update of atom (node) representation, $g$ and $f$ implemented by neural networks or mathematical operations, $d_{ij}$, $\theta_{ij}$ and $\phi_{ij}$ specify the 1-hop local neighborhood, $\tau_{ij}$ determines the orientation of the local neighborhood. We omit the final aggregation layer of ComENet to obtain the 3D representation corresponding to each atom $P_{\text{3D}}=\{\mathbf{v}_{i}\}_{i=1,2,\dots,N}$ rather than the entire molecular graph.
Subsequently, we input $P_{\text{3D}}$ along with the SMILES tokens embedding $T$ into the encoder. To facilitate a more comprehensive integration of SMILES and conformer, we aligned embeddings $T$ and $P_{\text{3D}}$ with atom index in the same vector space. We initially enlarge the length $P_{\text{3D}}$ from $N$ (number of atoms) to $M$ (number of tokens) by padding 0 to the indexes of non-atom tokens. In model input stage, we blend the $P_{\text{3D}}$ and $T$ to get the fusion embedding $F_{\text{3D}}$, allowing effective integration of 3D positional and sequential information at each atom position while preserving the alignment between them:
$F_{\text{3D}}=\lambda_{1}\times\operatorname{pad}(P_{\text{3D}})+\lambda_{2}\times T,$ (2)
where $\lambda_{1}$ and $\lambda_{2}$ are trainable parameters that can adaptively adjust weights of 3D position embedding $P_{\text{3D}}$ and SMILES tokens embedding $T$ during training. Besides, we still use Sinusoidal Position Embedding Vaswani et al. (2017) for tagging position of tokens.
4.2 Distance-weighted Attention
According to Li et al. (2017); Wan et al. (2022), multi-head attention calculation in vanilla transformer is insufficient and lacks chemical knowledge. Simultaneously, 3D distances between atoms within a molecule reflect the spatial relation between any pair of atoms in physical chemistry, such as bond energy, stability, electron affinity and interatomic force. Incorporating 3D distances into the multi-head attention calculation process enables the model to reschedule attention accordingly, which not only alleviates the redundancy in multi-head attention but also enables distance-based attention redistribution during the attention calculation.
Similar to the method of partitioning attention heads in Wan et al. (2022), we divide the multi-head attention into normal attention heads and spatial attention heads. Normal attention heads follow the traditional transformer manner to learn sequence-wide information, whereas spatial attention heads employ 3D distances as an external attention weight to enable attention redistributed based on distances. The whole module can be formulated as two steps: 1) Construct 3D distance weight, 2) Attention Redistribute & Weight Refine.
Construct 3D Distance Weight.
First, we construct a distance matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ from molecular conformer of molecule as:
$\mathbf{D}=\{\mathbf{d}_{i,j}\}_{i,j=1,2,\dots,N}\quad\mathbf{d}_{ij}=\|\mathbf{c}_{i}-\mathbf{c}_{j}\|_{2},$ (3)
where $N$ is number of atoms, $i,j$ are indexes of atoms pair, $\mathbf{c}_{i}$ is coordiante of $i$-th atom. To explore high-dimensional space, we employ Gaussian Basis Function Scholkopf et al. (1997) to transform the distance $\mathbf{d}_{ij}$ between each atom pair ($i$, $j$) into a higher dimension as:
$\psi_{(i,j)}^{k}=-\frac{1}{\sqrt{2\pi}\left|\sigma^{k}\right|}\exp\left(-\frac{1}{2}\left(\frac{\gamma_{(i,j)}\mathbf{d}_{ij}+\beta_{(i,j)}-\mu^{k}}{\left|\sigma^{k}\right|}\right)^{2}\right),$
where kk={1, …, K}, K is the number of Gaussian Kernel kernels. $\mu^{k}$ and $\sigma^{k}$ represent kernel center and scaling factor of $k$-th Gaussian Basis kernel. $\gamma_{(i,j)}$ and $\beta_{(i,j)}$ are learnable scalars related to bond type (e.g., SINGLE, DOUBLE) of atom pair ($i$, $j$). To obtain the 3D Distance weight $\Phi_{(i,j)}$, we use two layers of Non-Linear Layer with the activation function Gaussian Error Linear Units ($\operatorname{GELU}$) Hendrycks and Gimpel (2016) to encode $\psi_{(i,j)}$ as:
$\Phi_{(i,j)}=\mathbf{W_{2}}(\operatorname{GELU}(\mathbf{W_{1}}(\psi_{(i,j)}))),$ (5)
where $\psi_{(i,j)}=\operatorname{stack}([\psi_{(i,j)}^{1};\dots;\psi_{(i,j)}^{k}])$, $\mathbf{W_{1}},\mathbf{W_{2}}\in\mathbb{R}^{K\times K}$ are learnable parameters and 3D Distance weight $\Phi_{(i,j)}\in\mathbb{R}^{N\times N\times K}$.
Attention Redistribute & Weight Refine.
Second, we leverage the previously obtained 3D Distance weight $\psi_{(i,j)}$ to guide the model in redistributing attention weights by the internal spatial structure, instead of directly calculating attention scores between each token like vanilla Transformer. We divide the attention heads into two halves: normal attention head and spatial attention head. The normal attention head still captures the context of the SMILES sequence like original attention computation. The spatial attention head, on the other hand, focuses more on the molecular internal structure, limiting the receptive field of each atom and considering the 3D distance with other atoms. For the $i$-th token of the $l$-th encoder layer, the roll-out form of spatial attention head is:
$x_{i,\text{spatial}}^{l+1}$ $=\sum_{j\in N(i)}\operatorname{softmax}\left(\Phi_{(i,j)}^{l}\odot\frac{q_{i}k_{j}^{T}}{\sqrt{d}}\right)v_{j},$ (6)
${\left[q_{i},k_{j},v_{j}\right]}$ $=\left[h_{i}^{l}\mathbf{W^{Q}},h_{j}^{l}\mathbf{W^{K}},h_{j}^{l}\mathbf{W^{V}}\right],$
where $\Phi_{(i,j)}^{l}$ is almost the same as 3D Distance weight $\Phi_{(i,j)}$ but padding 0 in non-atom token for element-wise multiplication $\odot$, $\mathbf{W^{Q}}$, $\mathbf{W^{K}}$ and $\mathbf{W^{V}}$ are projection matrices for query $q$, key $k$, value $v$ in self-attention. Computed representations of normal attention head and spatial attention head are concatenated along the hidden dimension and then processed through a linear layer, forming the updated context for the next layer $h^{l+1}$, formally we have:
$h^{l+1}=\mathbf{W_{Linear}}\left(\left[x^{l+1}_{\text{normal}};x^{l+1}_{\text{spatial}}\right]\right).$ (7)
In addition, we perform Weight Refine to update our 3D Distance weight layer by layer. The Weight Refine module is a fully connected (FC) layer that takes concatenation of updated contexts of atoms pair ($i$, $j$) to refine the 3D attention weight among receptive field, which is formulated as:
$\Phi_{(i,j)}^{l+1}=\Phi_{(i,j)}^{l}+\operatorname{FC}\left(\left[h_{i}^{l+1};h_{j}^{l+1}\right]\right).$ (8)
4.3 SMILES Alignment
During the occurrence of chemical reactions, the majority of atoms remain unchanged Chen and Jung (2021); Zhong et al. (2022). We can utilize the atom-map to annotate the correspondence between atoms of product and reactants in reaction, which remains unchanged during the reaction. Furthermore, the correspondence of atoms can be readily transferred to the token correspondence in the SMILES sequence (see in Figure 1c) and construct SMILES Alignment Map (SAM) as:
$SAM_{ij}=\begin{cases}1&\text{ if }R_{i}\underset{\text{ mapped }}{\longleftrightarrow}P_{j}\\ 0&\text{ else }\end{cases}$ (9)
Inspired by Deshpande and Narasimhan (2020), we similarly employed the guided attention approach by introducing guidance loss $\mathcal{L}_{SA}$ to encourage alignment between the cross attention scores in the final layer of the decoder and SMILES Alignment Map, strengthening the attention weights between corresponding tokens.
Model Top-k Accuracy (%)
Reaction class known Reaction class unknown
1 3 5 10 1 3 5 10
Template-Based
RetroSim Coley et al. (2017) 52.9 73.8 81.2 88.1 37.3 54.7 63.3 74.1
GLN Dai et al. (2019) 64.2 79.1 85.2 90.0 52.5 69.0 75.6 83.7
LocalRetro Chen and Jung (2021) 63.9 86.8 92.4 96.3 53.4 77.5 85.9 92.4
Semi-Template-Based
RetroXpert Yan et al. (2020) 62.1 75.8 78.5 80.9 50.4 61.1 62.3 63.4
G2G Shi et al. (2020) 61.0 81.3 86.0 88.7 48.9 67.6 72.5 75.5
GraphRetro Somnath et al. (2020) 63.9 81.5 85.2 88.1 53.7 68.3 72.2 75.5
RetroPrime Wang et al. (2021) 64.8 81.6 85.0 86.9 51.4 70.8 74.0 76.1
MEGAN Sacha et al. (2021) 60.7 82.0 87.5 91.6 48.1 70.7 78.4 86.1
Graph2Edits Zhong et al. (2023) 67.1 87.5 91.5 93.6 55.1 77.3 83.4 89.4
Template-Free
SCROP Zheng et al. (2019) 59.0 74.8 78.1 81.1 43.7 60.0 65.2 68.7
Aug. Transformer Tetko et al. (2020) - - - - 48.3 - 73.4 77.4
GTA Seo et al. (2021) - - - - 51.1 67.6 74.8 81.6
Graph2SMILES Tu and Coley (2022) - - - - 52.9 66.5 70.0 72.9
Retroformer Wan et al. (2022) 64.0 82.5 86.7 90.2 53.2 71.1 76.6 82.1
NAG2G Yao et al. (2024) 67.2 86.4 90.5 93.8 55.1 76.9 83.4 89.9
Retro3D (Ours) 67.6 87.2 91.4 94.5 55.5 77.2 83.4 89.1
Table 1: Top-k accuracy for Retrosynthesis prediction on USPTO-50K. The metric values of other models in the table are taken from their original papers. Best performance is in bold.
4.4 Overall Loss
The overall loss is composed of the language modeling R-Drop lossWu et al. (2021) $\mathcal{L}_{LM}$ for Retrosynthesis, along with the SMILES Alignment guidance loss $\mathcal{L}_{SA}$.
The objectives of overall loss: i) approximate target in retrosynthesis, ii) ensure the robustness under different dropout, iii) strengthen the alignment relationship in cross-attention. The formula is as follows:
$\mathcal{L}=\mathcal{L}_{LM}^{(CE)}+\alpha\mathcal{L}_{LM}^{(KL)}+\beta\mathcal{L}_{SA},$ (10)
where $\mathcal{L}_{LM}^{(CE)}$ is Cross-Entropy loss to minimize prediction error against actual outputs, $\mathcal{L}_{LM}^{(KL)}$ is KL-divergence in R-Drop loss to strive for outputs of models with varying Dropout to be as similar as possible. $\mathcal{L}_{SA}$ is Cross-Entropy loss between cross-attention score of last decoder layer and SAM that we mention at Section 4.3. We set $\alpha$ to 0.5, $\beta$ to 1.0.
Model Top-k Validity (%) Top-k Round-Trip (%)
1 3 5 10 1 3 5 10
Graph2SMILES 99.4 90.9 84.9 74.9 76.7 56.0 46.4 34.9
RetroPrime 98.9 98.2 97.1 92.5 79.6 59.6 50.3 40.4
Retroformer 99.2 98.5 97.4 96.7 78.9 72.0 67.1 57.2
NAG2G 99.7 98.6 97.1 92.9 - - - -
Retro3D (Ours) 99.8 99.6 99.1 97.0 90.7 81.3 75.7 69.0
Table 2: Top-k SMILES Validity / Round-Trip Accuracy for Retrosynthesis prediction on USPTO-50K with reaction class unknown.
(a) (b) (c) Top-k Accuracy (%) Top-k Validity (%)
1 3 5 10 1 3 5 10
- - - 50.2 70.6 76.0 80.1 99.1 98.4 97.7 95.6
- - $\checkmark$ 52.8 73.8 80.7 86.4 99.3 98.9 98.5 95.3
$\checkmark$ - $\checkmark$ 53.3 73.1 79.1 84.6 99.6 99.2 98.4 94.9
- $\checkmark$ $\checkmark$ 54.1 73.2 79.1 85.2 99.7 99.4 98.6 94.2
$\checkmark$ $\checkmark$ - 54.1 73.0 80.0 86.5 99.7 99.6 98.9 95.9
$\checkmark$ $\checkmark$ $\checkmark$ 55.5 77.2 83.4 89.1 99.8 99.6 99.1 97.0
Table 3: Ablation study of Retro3D with reaction class unknown on USPTO-50K dataset. Module (a) is Atom-align Fusion; Module (b) is Distance-weighted Attention and Module (c) is SMILES Alignment;
5 Experiments
Datasets.
We use the popular Retrosynthesis benchmark dataset USPTO-50K [^1] USPTO-50K and USPTO-FULL Dataset can be obtained in https://github.com/Hanjun-Dai/GLN Schneider et al. (2016), which contains 50,016 atom-mapped reactions, to evaluate our model. The same data split as Coley et al. (2017) is applied to our experiments, resulting in 40,008, 5,001, and 5,007 reactions for the training, validation, and test sets, respectively. Conformers are generated using RDKit with the stochastic optimization algorithm Merck Molecular Force Field (MMFF). We then apply the same algorithm in Wan et al. (2022) to extract the ground truth SMILES Alignment Map. We also conduct experiments on the larger dataset USPTO-FULL, details can be seen in Appendix B.
Setting.
Our model is built on vanilla Transformer, consisting of 6 encoder layers and 6 decoder layers with 8 attention heads. The model dimension is set to 512. We employ the Adam optimizer Kingma and Ba (2014) with $(\beta_{1},\beta_{2})=(0.9,0.98)$ and train the model with batch size of 16. During training, we set an early-stop of 7 epochs under a maximum of 1000 epochs to obtain the 7 best models and average their parameters. Finally, the whole training process took approximately 30 hours on a single NVIDIA GeForce RTX 4090 GPU. Our code is available at https://anonymous.4open.science/r/Retro3D.
Evaluation.
During inference, we use beam search Tillmann and Ney (2003) decoding strategy with a beam size is 10. After that, we evaluate our model with Top-k Accuracy, which considers it correct only if Top-k predicted reactants of the beam search results are the same as those in the original test set. We verify Top-k Validity based on whether RDKit can recognize the Top-k predicted reactants. Top-k Round-Trip Accuracy which measures the percentage of predicted reactants that can lead back to the original product.
Baseline.
We compare our proposed method, Retro3D, with several strong baseline models from three types of retrosynthesis methods. We take RetroSim Coley et al. (2017), GLN Dai et al. (2019) and LocalRetro Chen and Jung (2021) to represent template-based methods. We take RetroXpert Yan et al. (2020), G2Gs Shi et al. (2020), GraphRetro Somnath et al. (2020), RetroPrime Wang et al. (2021), MEGAN Sacha et al. (2021) and Graph2Edits Zhong et al. (2023) to represent semi-template method. We take SCROP Zheng et al. (2019), Aug.Transformer Tetko et al. (2020), Graph2SMILES Zhong et al. (2023), GTA Seo et al. (2021) and Retroformer Wan et al. (2022), NAG2G Yao et al. (2024) to represent template-free method. We did not choose pre-trained methods for comparison as they are trained on larger datasets. We apply Root-align data augmentation Zhong et al. (2022) and conducted experiments on the proposed model Retro3D, choosing on-the-fly augmentation instead of directly expanding the training dataset.
5.1 Performance
Top-k Accuracy.
As shown in Table 1, when the reaction class is known, Retro3D achieves Top1-10 results that outperform all template-free/semi-template baselines and most state-of-the-art template-based baselines. Notably, our model exceeds both template-based and semi-template methods, despite their reliance on predefined templates or RDKit editing. When the reaction class is unknown, our model still achieves improvements in Top1-3 accuracy within the template-free category. Achieving high-confidence predictions (i.e., Top-1 acc), which narrows the search space for synthesis routes and enhances efficiency, makes our model particularly well-suited for real-world scenarios. Performance across each reaction classes in USPTO-50K dataset are provided in Appendix A. Result on USPTO-FULL dataset can be seen in Appendix B.
(a) Polychiral
(b) Heteroaromatic
(c) Fused / Bridged Rings
(d) Complex
Figure 2: Case study for intricate structures. Each subfigure shows predictions of different methods: Retro3D (Top), Transformer (Bottom).
Top-k Validity.
We take three typical template-free methods Graph2SMILES, RetroPrime, Retroformer and NAG2G as baseline models to verify the predicted SMILES validity of our model. We refrained from selecting template-based and semi-template methods, as SMILES constructed based on template information ensure validity. In contrast, template-free methods often struggle to generate valid SMILES without extra chemical guidance. As shown in Table 2, our models are more capable of generating valid SMILES than other baseline models. Especially for Top5-10 reactants, which are far ahead of other models. This indicates that incorporating 3D conformer information can assist the model in grasping the structures involved in chemical reactions, thus preventing the generation of SMILES that violate chemical principles.
Top-k Round-Trip Accuracy.
To assess the accuracy of our predicted synthesis plans, we use the Molecule Transformer Schwaller et al. (2019) as the benchmark reaction prediction model and calculate the Top-k Round-Trip Accuracy. The results in Table 2 clearly show that our model significantly outperforms other baselines, achieving a remarkable 12% improvement in Top-1 Round-Trip Accuracy compared to the best baseline.
5.2 Ablation Study
To explore the impact of three modules, we design an ablation study under the settings of unknown reaction class. The results are shown in Table 3 in the form of top-k accuracy and demonstrate necessity of each module for Retro3D. ② highlights the necessity of module (c) SMILES Alignment. The improvement in top-1 accuracy from ③ to ② suggests that the model is capable of using 3D positional information to infer the correct results with higher confidence after introducing module (a) Atom-align Fusion. Notably, the declines in top3-10 accuracy is speculated on the reliance upon fixed 3D position embedding, which constrains the search space of model. Comparing ④ and ②, it can be seen that employing (b) Distance-weighted Attention improves performance in all top-k accuracy, even for top3-10. We conjecture that redistributing attention enables the model to focus on internal structure of molecules, while Gaussian Kernel Function grants the model a larger search space. The results of ⑥ show that the combination of (a), (b) and (c) has achieved the best performance of model, enabling it to focus on internal spatial structures while maintaining an extensive search space based on comprehensive 3D positional information.
5.3 Case Study
To verify that the proposed method Retro3D can handle molecules with more intricate spatial structures, we chose 4 types of products as representatives:
  • Polychiral: a product with two or more chiral centers (’@/@@’ in SMILES), leading to complex stereochemistry and multiple isomers.
  • Heteroaromatic: a product containing at least one heteroatom as part of its aromatic ring structure.
  • Fused / Bridged Rings: product features multiple ring structures joined together, either sharing common atoms or connected by bridge atoms.
  • Complex: a product encompassing two or more features mentioned above.
We conducted experiments comparing Retro3D with a Vanilla Transformer, analyzing Top1-3 predictions, as illustrated in Figure 2. The rows in each subfigure represent the results of Retro3D and Transformer, respectively, with Ground Truth and invalid SMILES highlighted in bold red.
The results show that when dealing with molecules of intricate spatial structures, Retro3D can infer reasonable reactants with high confidence (often at top1, top2), while Transformer without 3D information struggles to generate accurate predictions. Notably, for Complex products, Retro3D successfully deduces the molecular skeleton and achieves the Ground Truth at Top-1, while the Vanilla Transformer fails to produce any valid SMILES within the Top-3 predictions.
6 Conclusion
We propose Retro3D, an end-to-end Transformer that integrates 3D conformer for Retrosynthesis. With the proposed Atom-align Fusion module, we can adaptively integrate token and 3D position information ensuring the alignment between them. Next, we propose Distance-weighted Attention mechanism to guide and refine the redistribution of self-attention calculation with spatial distances. Our model achieves a new state-of-the-art in template-free methods and is also highly competitive with template-based and semi-template methods. Furthermore, to validate the practicality of the model, we predicted synthesis routes for various candidate drug molecules, demonstrating its potential to revolutionize the prediction of complex chemical synthesis processes in future synthesis route design tasks. The results are provided in Appendix G.
Appendix A USPTO-50K Dataset
A.1 Performance across Each Reaction Class
In Table 3, we present the top-$k$ accuracy (kk=1, 3, 5, 10) of Retro3D on the USPTO-50k dataset, where the reaction types are unknown during the training process. The results are categorized by the ground truth reaction types.
Our method seems to predict certain reaction types more accurately than other methods, especially those that are more concerned with 3D spatial structure, such as in Heterocycle formation, where the 3D structure of the molecule directly affects the stereoselectivity and yield of the cyclization reaction. It is challenging for other methods to deduce accurate results without 3D structural information.
Reaction Class Reaction Fraction (%) Top-$k$ Accuracy (%)
1 3 5 10
Heteroatom alkylation and arylation 30.3 56.6 80.4 86.0 92.0
Acylation and related processes 23.8 68.5 89.9 94.3 97.7
Deprotections 16.5 54.5 78.5 85.9 91.9
C-C bond formation 11.3 41.6 63.1 73.9 83.4
Reductions 9.2 56.4 78.6 85.3 90.8
Functional group interconversion 3.7 39.4 56.4 64.6 76.1
Heterocycle formation 1.8 47.0 70.7 72.9 81.8
Oxidations 1.6 61.2 82.8 91.5 95.3
Protections 1.3 68.1 80.5 89.4 93.9
Functional group addition 0.5 80.2 85.2 89.6 98.4
Table 4: The top-$k$ accuracy of each reaction class on USPTO-50k dataset, when trained with reaction class unknown.
A.2 Performance on Different Conformer Generator
To explore the impact of different conformer generators on model performance, we conduct an ablation study. As shown in Table 5, in addition to the default RDkit generator, we evaluate the performance of the chemical simulation software Schrödinger, which is widely regarded as capable of producing higher-quality conformations. The results show that the overall performance remains comparable. This observation suggests that our model exhibits strong robustness, as the quality of conformations has minimal impact on the final performance of Retro3D.
Conformer Generator Top-k Accuracy (%)
1 3 5 10
Schrödinger Watts et al. [2010] 55.2 77.1 83.8 90.1
Default (RDkit) 55.5 77.2 83.4 89.1
Table 5: The performance with different conformer generator on USPTO-50k with reaction class unknown.
A.3 Performance on Different 3D Representation
This section explores the impact of 3D representations on model performance. Table 6 shows that as the 3D representation capacity increases (from top to bottom), the model’s performance improves, indicating a positive correlation between them.
3D Representation Network Top-k Accuracy (%)
1 3 5 10
SchNet Schütt et al. [2018] 51.0 69.5 73.5 76.3
SphereNet Coors and others [2018] 53.1 77.1 83.1 87.8
Default (ComENet) 55.5 77.2 83.4 89.1
Table 6: The performance with different conformer generator on USPTO-50k with reaction class unknown.
Appendix B USPTO-FULL Datasets
We also use a more extensive collection, USPTO-FULL Dataset, which contains 1,013,118 atom-mapped reactions without any reaction class information. By experimenting with it, we can evaluate the proposed method’s performance, generalization potential, and scalability, offering valuable insights into its practical application in retrosynthesis. Additionally, the dataset includes samples with more intricate structures, which allows for further validation of our model’s capability to handle molecules with intricate spatial configurations.
B.1 Preprocess
The USPTO-FULL raw data contains errors and other issues, so we implemented the following steps to clean and refine the dataset:
  • Exclude reactions with invalid SMILES, including those with empty SMILES strings.
  • Exclude reactions where the product contains atoms not present in the reactants.
  • Exclude reactants if none of their atoms are found in the product.
  • Exclude reactions where multiple atoms share the same atom map number.
  • Exclude reactions where the product is composed of less than five atoms.
The final version shows an approximate 4% reduction compared to the raw data. The dataset is divided into training, validation, and test sets with sizes of 768K, 96K, and 96K, respectively.
B.2 Performance
Table 5 provides a performance comparison of various models on the USPTO-Full dataset. The findings reveal that Retro3D consistently outperforms previous baselines across all metrics. Furthermore, it is observed that template-based/semi-template methods, despite their strong performance on the USPTO-50k dataset, struggle when applied to the USPTO-Full dataset. This suggests that these methods face challenges in scaling due to their reliance on pre-defined rules. In contrast, our approach shows better performance when handling large-scale datasets and molecules with complex spatial structures.
Model Top-k Accuracy (%)
1 3 5 10
Template-Based
RetroSim 32.8 - - 56.1
GLN 35.8 - - 60.8
LocalRetro 39.3 - - 63.7
Semi-Template-Based
RetroPrime 44.1 59.1 62.8 68.5
MEGAN 33.6 - - 63.9
Graph2Edits 44.0 60.9 66.8 72.5
Template-Free
Aug. Transformer 46.2 - - 73.3
GTA 46.6 - - 70.4
Graph2SMILES 45.7 - - 63.4
NAG2G 49.7 64.6 69.3 74.0
Retro3D (Ours) 50.8 67.5 72.9 78.0
Table 7: Top-k accuracy for Retrosynthesis prediction on USPTO-FULL without reaction class. The metric values of other models in the table are taken from their original papers. Best performance is in bold.
Appendix C Implementation Details
C.1 3D Distance Matrix Construction
In the Distance-weighted Attention mechanism, a 3D Distance Matrix needs to be constructed first. Algorithm 1 is the pseudo code for the construction of the 3D Distance Matrix $\mathbf{D}$ from a molecular canonical-SMILES $S$.
The algorithm begins by converting the SMILES into a molecular object $M$, followed by generating a conformer for the molecule and obtaining the coordinates of all atoms within it. Subsequently, an empty distance matrix is initialized. For each pair of atom tokens derived from tokenizing the SMILES, if both tokens represent atoms, the algorithm calculates the Euclidean distance between these two atoms and assigns this value to the corresponding position in the distance matrix; if they are not atom tokens, the algorithm sets the corresponding matrix element to 0. Finally, the algorithm returns the constructed 3D distance matrix $\mathbf{D}$.
Algorithm 1 3D Distance Matrix Construction
Input: molecule canonical SMILES $S$
Output: 3D Distance Matrix $\mathbf{D}$
1: Let $M=\operatorname{MolFromSmiles}(S)$.
2: Generate conformer by $\operatorname{MMFFOptimizeMolecule}(M)$.
3: Get coordinates of atoms $\mathbf{c}=\operatorname{GetPositions}(M)$.
4: Initialize $\mathbf{D}$ as an empty array.
5:for $t_{i},t_{j}\in\operatorname{SmiTokenizer}(S)$ do
6:  if $t_{i}$, $t_{j}$ are atom tokens then
7:   $\mathbf{D_{i,j}}=\operatorname{norm}(\mathbf{c_{i}}-\mathbf{c_{j}})$
8:  else
9:   $\mathbf{D_{i,j}}=0$
10:  end if
11:end for
12:return $\mathbf{D}$
C.2 Root-align SMILES Construction
Algorithm 2 is the pseudo code for the construction of the reactants Root-align SMILES $S_{R}$(+R) from a reactants canonical SMILES $S$.
First, the algorithm extracts the atom-mapped number from the product SMILES $S_{P}$, then iterates through the SMILES for each reactant, looking for the same atom-mapped number as in the $S_{P}$. Once a match is found, the algorithm uses the RDkit toolkit to generate a fixed-root SMILES centered on the corresponding atom and returns the constructed Root-align SMILES $S_{R}$(+R).
Algorithm 2 Root-align SMILES Construction
Input: atom-mapped product SMILES $S_{P}$ and atom-mapped reactants SMILES $S_{R}$
Output: reactants R-SMILES $S_{R}$(+R)
1: Get atom-mapped number list $am(S_{P})$ of $S_{P}$ by Regular Expression Matching.
2:for $pNum\in am(S_{P})$ do
3:  for $r_{i}\in\operatorname{SmiTokenizer}(S_{R})$ do
4:   if $r_{i}$ is atom-mapped tokens and $am(r_{i})==pNum$ then
5:    Make $r_{i}$ as root-atom of $S_{R}$ and get reactants R-SMILES $S_{R}$(+R)
6:    return $S_{R}$(+R)
7:   end if
8:  end for
9:end for
10:return $S_{R}$(+R)
C.3 SMILES Alignment Map Construction
The ground truth SMILES Alignment Map $SAM$ between the product SMILES $S_{P}$ and the reactants SMILES $S_{R}$ is computed as Algorithm 3.
The algorithm first initializes an empty $SAM$ to store the correspondence between $S_{R}$ and $S_{P}$. For each atom-mapped token in $S_{R}$, the algorithm looks for a corresponding token with the same atom-mapped number in $S_{P}$. Once this correspondence is found, the algorithm adds the pair of tokens to the $SAM$. This process will traverse forward and backward until there are no more matching tokens. Finally, the algorithm returns the constructed SMILES Alignment Map $SAM$.
Algorithm 3 SMILES Alignment Map Construction
Input: atom-mapped product SMILES $S_{P}$ and atom-mapped reactants SMILES $S_{R}$
Output: SMILES Alignment Map SAMSAM
1: Initialize SMILES Alignment Map $SAM$.
2:for $r_{i}\in\operatorname{SmiTokenizer}(S_{R})$ do
3:  if $r_{i}$ is not visited and $r_{i}$ is an atom-mapped token then
4:   Find the token $p_{j}\in\operatorname{SmiTokenizer}(S_{p})$ with the same atom-map number as $r_{i}$: ${am}(p_{j})=={am}(r_{i})$.
5:   while $r_{i}==p_{j}$ or ${am}(p_{j})=={am}(r_{i})$ do
6:    Add token alignment pair $r_{i}:p_{j}$} into SAMSAM.
7:    $i=i+1$ and $j=j+1$.
8:   end while
9:  end if
10:end for
11:for $r_{i}:p_{j}$} $\in SAM$ do
12:  i=i1i=i-1 and j=j1j=j-1.
13:  while $r_{i}==p_{j}$ and $r_{i}$ is not an atom token do
14:   Add token alignment pair $r_{i}:p_{j}$} into SAMSAM.
15:   i=i1i=i-1 and j=j1j=j-1
16:  end while
17:end for
18:return SAMSAM
Appendix D Hyperparameters Setting
Hyperparameter value
Layers 6
Attention heads 8
Embedding dim 512
FFN hidden dim 2048
Gaussian kernel channels 512
FFN dropout 0.1
Attention dropout 0.1
Embedding dropout 0.1
Coefficient for KL-loss ($\alpha$) 0.5
Activation function ReLU
Vocabulary size 84
Batch size 16
Max training epoch 500
Beam Search Size 10
Warmup steps 8000
Weight decay 1e-3
Learning Rate Scheduler Noam
Learning rate decay 0.001
Peak learning rate 23.8
Adams $\epsilon$ 0.5
Adams ($\beta_{1},\beta_{2}$) (0.9, 0.98)
Table 8: Retro3D hyperparameters setup during training
Appendix E Visualization
E.1 Self-Attention Map
For better interpretability, we visualize the molecular self-attention maps and distance pair matrices as shown in Figure 3. It is evident that the sum of the spatial attention heads (Attention Heads 5, 6, 7, 8 in Figure 3) closely resembles the distance matrix, suggesting that spatial information has been integrated into the self-attention computation. However, when we check these heads independently, different patterns emerge. For example, the normal attention head weights are asymmetric, focusing more on self-token interactions, whereas the distance matrix tends to interact with other tokens. Additionally, long-range interactions can be captured through spatial attention (such as in Attention Heads 7 and 8).
Figure 3: Visualization on self-attention map of multi heads independently.
E.2 Cross-Attention Map
SMILES Alignment between product and reactants and Cross Attention Score on first head, respectively. The left most figure is the ground truth alignment matrix. It is easy to find that cross-attention captures the corresponding information of the SMILES Alignment Map very well and assists the model generation process in the Decoder stage, which also promotes the final excellent performance of our model.
Figure 4: Visualization on cross-attention map on SMILES Alignment.
Appendix F Case Study
Here we show statistic analysis and more illustrations in Case Study on Intricate Molecules.
Type Model Top-k Accuracy (%) Top-k Validity (%)
1 3 5 10 1 3 5 10
Polychiral Retro3D 42.5 65.4 71.5 79.2 98.6 98.8 98.6 97.7
Trans. 38.5 59.0 67.5 73.9 97.3 97.3 96.2 92.6
Heteroaromatic Retro3D 55.3 77.1 84.0 89.9 99.3 98.9 98.7 96.1
Trans. 51.2 71.9 79.0 84.7 98.2 97.6 96.4 91.1
Fused / Bridged Rings Retro3D 55.4 76.4 83.8 88.8 99.4 99.1 98.7 96.7
Trans. 50.8 71.1 77.8 83.3 99.2 98.1 97.1 91.6
Complex Retro3D 54.3 75.6 82.5 88.3 99.4 99.0 98.7 96.2
Trans. 49.5 70.3 77.5 83.2 98.3 97.6 96.3 92.0
Table 9: Performance comparison between Retro3D and Transformer models across four types of molecular structures.
Figure 5: More Case
Appendix G Synthesis route
By applying our product-reactant variant recursively, we verify our method with several multistep retrosynthesis examples reported in the literature, including febuxostat Cao et al. [2016], salmeterol Guo et al. [2011], an allosteric activator for GPX4 Lin et al. [2020], a 5-HT6 receptor ligand Nirogi et al. [2015], Nirmatrelvir Hammond and others [2022] and Osimertinib Finlay and others [2014]. As shown in Fig. 6 and Fig. 7, our method successfully predicts the complete synthetic pathway for these examples.
Figure 6: Multistep retrosynthesis predictions by our method. (a) Febuxostat. (b) Salmeterol. (c) Allosteric Activator for GPX4. (d) 5-HTe Receptor Ligand.
Figure 7: Multistep retrosynthesis predictions by our method. (e) Nirmatrelvir. (f) Osimerinib.